# A Novel Framework for Uncertainty-Driven Adaptive Exploration

### v5

@@ 1. Introduction
To test our framework, we create different instances of ADEU, each corresponding to a different uncertainty-measuring mechanism. We test our method in the most complex/difficult among the robotic tasks in towers2024gymnasium and in the DeepSea domain—a domain which tests the agent’s ability for deep exploration. Our experiments demonstrate that the performance of an ADEU instance depends on the ability of each uncertainty-measuring mechanism to quantify uncertainty. However, in our experiments, the instance of ADEU employing a specific uncertainty-measuring mechanism, outperforms the original exploration method that employs the same uncertainty-measuring mechanism. Moreover, there is always some instance of ADEU that performs better than all of its opponents. As such, ADEU can be considered as a ‘plug-and-play’ adaptive exploration framework, that can be coupled with any uncertainty-measuring mechanism; and exhibits the ability to improve the performance of DRL algorithms that utilize the corresponding uncertainty-measuring mechanism in their exploration process.
@@ 2.1. Formal Framework Description
Theorem 4 shows that an ADEU agent, when provided with an effective uncertainty-measuring mechanism, will solve an Increasing-Reward Single-Agent Game. The theorem can be extended to guarantee efficient (in expectation) ADEU behavior in environments with stochastic reward or transition functions also. ADEU’s efficiency arises from the fact that it does not needlessly explore, but instead replicates the parts of a “known” or “learnt” trajectory that it trusts. But this trajectory might be sub-optimal. To keep improving the learnt trajectory, ADEU also performs the so-called rollout episodes. In more detail, at the beginning of every episode, with probability $\rho$, ADEU performs exploration using as uncertainty mechanism the $f_{\text{rollout}}(s)=c$, where $c$ is a constant number. Hence, for that specific “rollout” episode, ADEU takes the chance to explore so as to improve the learnt trajectory. We should note that in the case of a well-constructed $f(s)$ that smoothly decreases, $\rho$ should and can simply be set to 0. For simplicity, in our implementation, we treat $\rho$ as a constant that does not change throughout the training process. Also, we set $f_{\text{rollout}}(s)=c$ to present a simple version of our framework, but, in more complex environments, this can change too.
@@ 2.2. Generality of the ADEU framework
Given that ADEU can subsume existing adaptive exploration methods as special cases, it is reasonable to expect that it can replicate their performance in the tested domains. Thus, ADEU is well-suited for tasks with both dense and sparse reward structures. For instance, ADEU can solve the extremely demanding DeepSea domain osband2019deepSea by emulating the dabney2021ez_greedy adaptive exploration method.
@@
Figure 3 (top subfigure) shows both the actual policy $\pi(s)$ and the action selected from the ADEU agent (so, using Equation 1) during a single episode in our Frozen Lake environment. In addition, the bottom subfigure shows the value of $g(f(s))$ during the same episode. As expected, at the beginning of the episode, $g(f(s))$ outputs low values since the agent visits states that it is certain about. Hence, the action selected by the ADEU agent coincides with that dictated by $\pi(s)$. As the agent visits states with high uncertainty, it does not trust its so far learned policy and starts exploring them. As of that, the action selected by ADEU and the policy $\pi(s)$ differ at the end of the episode, since $g(f(s))$ increases. Note also that the action suggested by $\pi(s)$ in the later timesteps of the episode is not correct. However, the action sampled for exploration was the correct one, allowing the agent to “survive” for more timesteps and gain useful experience. Thus, indeed, the exploration phase should have begun at the later timesteps; had it started earlier, the agent would not have “survived” to extend the known optimal policy.  [^7] We expect the same behaviour in more difficult multi-dimensional domains such as the ones described in Section 3. However, we cannot create a figure similar to Figure 3 for those, since both $\pi(s),a\in\mathbb{R}^{n}$, where $n>2$.
@@ 2.3. ADEU for safe exploration
where $f_{\text{safety}}(s)$ is the expectation of the environment’s cost function $\mathbb{E}_{s,a}\max_{a}\{g_{\text{cost}}(s,a)\}$ (where $g_{\text{cost}}(s,a)$ follows the standard definition described in altman1999cmdp), which is learned by interactions.
@@
In Eq. 6 $f_{\text{uncertainty}}(s)$ is the aforementioned uncertainty-measuring mechanism and all other notation (except $\lambda$) follows Eq. 5. Note that $\lambda$ generally acts as a term (or mechanism) that ensures the safety term “dominates” when safety is critical. For example, $\lambda$ can take “conditioned” values for a state $s$ e.g., $\lambda=[max_{a}\{Q(s,a)\}$ if $\mathbb{E}_{s,a}\max_{a}\{g_{\text{cost}}(s,a)\}>max_{a}\{Q(s,a)\}$, otherwise 1 ]]. Therefore, the $f_{\text{safety}}(s)$ term will be amplified when the expected cost is high.
@@ 3. Experimental Evaluation
We conduct our experiments in standard MuJoCo domains towers2024gymnasium. Specifically, we used the five domains considered the most difficult according to the official documentation. [^10] https://gymnasium.farama.org/environments/mujoco/ . We did not use gymnasium_robotics2023github since its domains do not have the “early termination” condition. Such domains do not satisfy Definition 1. However, the ADEU agents still achieve (as later described) superior performance against their opponents. In these domains, the agent should explore the continuous multidimensional action space to find the best policy. A dense reward function is used to describe these robotic tasks. However, ADEU can also be used in environments with a sparse reward function, when using appropriate uncertainty-measuring mechanisms e.g., Badia2020NeverGiveUp; nikolov2018informationdirected.
@@
To conclude the experimental section, it is clear that the ADEU framework, paired with efficient uncertainty-measuring mechanisms, can be utilized for both demanding control tasks and domains requiring deep exploration.

### v6

@@ 1. Introduction
To test our framework, we create different instances of ADEU, each corresponding to a different uncertainty-measuring mechanism. We test our method in the most complex/difficult among the robotic tasks in towers2024gymnasium, and in the DeepSea domain. Our experiments demonstrate that the performance of an ADEU instance depends on the ability of each uncertainty-measuring mechanism to quantify uncertainty. However, in our experiments, the instance of ADEU employing a specific uncertainty-measuring mechanism, outperforms the original exploration method that employs the same uncertainty-measuring mechanism. Moreover, there is always some instance of ADEU that performs better than all of its opponents. As such, ADEU can be considered as a ‘plug-and-play’ adaptive exploration framework, that can be coupled with any uncertainty-measuring mechanism; and exhibits the ability to improve the performance of DRL algorithms that utilize the corresponding uncertainty-measuring mechanism in their exploration process.
@@ 2.1. Formal Framework Description
Proof.
The proof can be found in the Appendix. ∎
Theorem 4 shows that an ADEU agent, when provided with an effective uncertainty-measuring mechanism, will solve an Increasing-Reward Single-Agent Game. The theorem can be extended to guarantee efficient (in expectation) ADEU behavior in environments with stochastic reward or transition functions also. ADEU’s efficiency arises from the fact that it does not needlessly explore, but instead replicates the parts of a “known” or “learnt” trajectory that it trusts. But this trajectory might be sub-optimal. To keep improving the learnt trajectory, ADEU also performs the so-called rollout episodes. In more detail, at the beginning of every episode, with probability $\rho$, ADEU performs exploration using as uncertainty mechanism the $f_{\text{rollout}}(s)=c$, where $c$ is a constant number. Hence, for that specific “rollout” episode, ADEU takes the chance to explore so as to improve the learnt trajectory. We should note that in the case of a well-constructed $f(s)$ that smoothly decreases, $\rho$ should and can simply be set to 0. For simplicity, in our implementation we treat $\rho$ as a constant that does not change throughout training. Also, we set $f_{\text{rollout}}(s)=c$ to present a simple version of our framework, but, in more complex environments, this can change too.
@@ 2.2. Generality of the ADEU framework
Given that ADEU can subsume existing adaptive exploration methods as special cases, it is reasonable to expect that it can replicate their performance in the tested domains. Thus, ADEU is well-suited for tasks with both dense and sparse reward structures. For instance, ADEU solves the extremely demanding DeepSea domain osband2019deepSea by emulating Dabney et al.’s dabney2021ez_greedy adaptive exploration method.
@@
Figure 3 (top subfigure) shows both the actual policy $\pi(s)$ and the action selected from the ADEU agent (so, using Equation 1) during a single episode in our Frozen Lake environment. In addition, the bottom subfigure shows the value of $g(f(s))$ during the same episode. As expected, at the beginning of the episode, $g(f(s))$ outputs low values since the agent visits states that it is certain about. Hence, the action selected by the ADEU agent coincides with that dictated by $\pi(s)$. As the agent visits states with high uncertainty, it does not trust its so far learned policy and starts exploring them. As of that, the action selected by ADEU and the policy $\pi(s)$ differ at the end of the episode, since $g(f(s))$ increases. Note also that the action suggested by $\pi(s)$ in the later timesteps of the episode is not correct. However, the action sampled for exploration was the correct one, allowing the agent to “survive” for more timesteps and gain useful experience. Thus, indeed, the exploration phase should have begun at the later timesteps; had it started earlier, the agent would not have “survived” to extend the known optimal policy.  [^7] We expect the same behaviour in more difficult multi-dimensional domains such as the ones described in Section 3. However, we cannot create a figure similar to Figure 3 for those, since both $\pi(s),a\in\mathsf{R}^{n}$, where $n>2$.
@@ 2.3. ADEU for safe exploration
where $f_{\text{safety}}(s)$ is the expectation of the environment’s cost function $\mathsf{E}_{s,a}\max_{a}\{g_{\text{cost}}(s,a)\}$ (where $g_{\text{cost}}(s,a)$ follows the standard definition described in altman1999cmdp), which is learned by interactions.
@@
In Eq. 6 $f_{\text{uncertainty}}(s)$ is the aforementioned uncertainty-measuring mechanism and all other notation (except $\lambda$) follows Eq. 5. Note that $\lambda$ generally acts as a term (or mechanism) that ensures the safety term “dominates” when safety is critical. For example, $\lambda$ can take “conditioned” values for a state $s$ e.g., $\lambda=[max_{a}\{Q(s,a)\}$ if $\mathsf{E}_{s,a}\max_{a}\{g_{\text{cost}}(s,a)\}>max_{a}\{Q(s,a)\}$, otherwise 1 ]]. Therefore, the $f_{\text{safety}}(s)$ term will be amplified when the expected cost is high.
@@ 3. Experimental Evaluation
We first conducted experiments in standard MuJoCo domains towers2024gymnasium. Specifically, we used the five domains considered the most difficult according to the official documentation. [^10] https://gymnasium.farama.org/environments/mujoco/ . We did not use gymnasium_robotics2023github since its domains do not have the “early termination” condition. Such domains do not satisfy Definition 1. However, the ADEU agents still achieve (as later described) superior performance against their opponents. In these domains, the agent should explore the continuous multidimensional action space to find the best policy. A dense reward function is used to describe these robotic tasks. However, ADEU can also be used in environments with a sparse reward function, when using appropriate uncertainty-measuring mechanisms e.g., Badia2020NeverGiveUp; nikolov2018informationdirected.
@@
It is clear from our experiments that ADEU, paired with efficient uncertainty-measuring mechanisms, can be utilized for both demanding control tasks and domains requiring deep exploration.
@@ 5. Conclusions and Future Work
{acks}
The research described in this paper was carried out within the framework of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union - NextGenerationEU (Implementation Body: HFRI. Project name: DEEP-REBAYES. HFRI Project Number 15430).